OpenStudy (221emily):

I am having a hard time working through this problem. I do not wish to be told the answer! I would just like some help in seeing if I did the problem correctly :)

OpenStudy (221emily):

What is an equation in slope-intercept form for the line that passes through the point (1,-3) and (3,1)? A.) Y=3x+1 B.) Y=x-3 C.) Y=2x+5 D.) Y=2x-5 my answer is D.

OpenStudy (221emily):

omg sissssssssssssssss lol

OpenStudy (anonymous):

lol i took that last year and i jut plugged in the number and a hich one made ene

OpenStudy (221emily):

oh haha ok I still think that is D